Determine the voltage output of the circut


1
Expert's answer
2021-05-20T10:10:39-0400

"R_1=2\ \Omega,\ R_2=3\ \Omega,\ I=5\ A"

Output voltage:

"U=I(R_1+R_2)=5(2+3)=25\ V"